OpenStudy (anonymous):

@satellite73 okay so the questions is "find all points x,y on the curve where the line tangent to the curve has slope -1/2. so the curve given is y=(2x-5)/(x^2-4) and the derivative was [-2(x-4)(x-1)]/[(x^2-4)^2]

OpenStudy (anonymous):

ah in that case, set the derivative equal to -1/2 and solve for x

OpenStudy (anonymous):

yeah thats the hard part. I don't know how. I tried doing it by calculator thru intersection but it wouldnt work

OpenStudy (anonymous):

ok we can do that part

OpenStudy (anonymous):

okay so we have 2(x-4)(x-1)]/[(x^2-4)^2=-1/2

OpenStudy (anonymous):

i think you might be off by a minus sign. let me check carefully. \[\frac{-2(x-4)(x-1)}{(x^2-4)^2}=-\frac{1}{2}\]

OpenStudy (anonymous):

oh no you have it on the first line. ok

OpenStudy (anonymous):

so the first step would be to get rid of the denominator right?

OpenStudy (anonymous):

this means \[\frac{2(x-4)(x-1)}{(x^2-4)^2}=\frac{1}{2}\] so \[4(x-4)(x-1)=(x^2-4)^2\]

OpenStudy (anonymous):

now that i have started i have no clue as to how you are supposed to continue because you have a polynomial of degree 4

OpenStudy (anonymous):

i got this from wolfram http://www.wolframalpha.com/input/?i=4%28x-4%29%28x-1%29%3D%28x%5E2-4%29%5E2 two real solutions, but how you are supposed to get the second one i have absolutely no idea

OpenStudy (anonymous):

well thanks anyway. I'll just ask one of the math teachers tomorrow early morning.

OpenStudy (anonymous):

i checked and your derivative is correct. maybe they gave you this problem without thinking it through